OpenStudy (anonymous):

what is the inverse of (x+2)^3 Need ASAP Will medal

OpenStudy (anonymous):

solve \[x=(y+2)^3\]for \(y\) in two steps

OpenStudy (anonymous):

a) take the cubed root of both sides b) subtract 2 from both sides

OpenStudy (solomonzelman):

In general (for odd integer \(n\)). I added a b coefficient, or else it would have been too explicit, although it already is pretty simple. Suppose you want to find the inverse of the following function, \(\color{#000000 }{ \displaystyle f(x)=b(x+a)^n }\) ■ replacing f(x), by y for convinience \(\color{#000000 }{ \displaystyle y=b(x+a)^n }\) ■ swiping the x and the y \(\color{#000000 }{ \displaystyle x=b(y+a)^n }\) ■ swiping the x and the y \(\color{#000000 }{ \displaystyle x=b(y+a)^n }\) ■ solve for y. \(\color{#000000 }{ \displaystyle \sqrt[n]{x}=\sqrt[n]{b(y+a)^n} }\) \(\color{#000000 }{ \displaystyle \sqrt[n]{x}=\sqrt[n]{~b~}\cdot \sqrt[n]{(y+a)^n} }\) \(\color{#000000 }{ \displaystyle \sqrt[n]{x}=\sqrt[n]{~b~}(y+a) }\) \(\color{#000000 }{ \displaystyle \sqrt[n]{x}=y\sqrt[n]{~b~}+a\sqrt[n]{~b~} }\) \(\color{#000000 }{ \displaystyle \sqrt[n]{x}-a\sqrt[n]{~b~}=y\sqrt[n]{~b~} }\) \(\color{#000000 }{ \displaystyle \frac{\sqrt[n]{x}-a\sqrt[n]{~b~}}{\sqrt[n]{~b~}}=y }\) ■ since y is an inverse function, it has a notation \(\). \(\color{#000000 }{ \displaystyle f^{-1}(x)=\frac{\sqrt[n]{x}-a\sqrt[n]{~b~}}{\sqrt[n]{~b~}} }\)